### v1

# Hypergrah-Enhanced Dual Convolutional Network for Bundle Recommendation

Abstract
Bundle recommendations strive to offer users a set of items as a package named bundle, enhancing convenience and contributing to the seller’s revenue. While previous approaches have demonstrated notable performance, we argue that they may compromise the ternary relationship among users, items, and bundles. This compromise can result in information loss, ultimately impacting the overall model performance. To address this gap, we develop a unified model for bundle recommendation, termed hypergraph-enhanced dual convolutional neural network (HED). Our approach is characterized by two key aspects. Firstly, we construct a complete hypergraph to capture interaction dynamics among users, items, and bundles. Secondly, we incorporate U-B interaction information to enhance the information representation derived from users and bundle embedding vectors. Extensive experimental results on the Youshu and Netease datasets have demonstrated that HED surpasses state-of-the-art baselines, proving its effectiveness. In addition, various ablation studies and sensitivity analyses revealed the working mechanism and proved our effectiveness. Codes and datasets are available at https://github.com/AAI-Lab/HED
Index Terms:
Bundle Recommendation, Graph Convolutional Network, Hypergraph, Recommender System
I Introduction
Recommender system is a type of system that automatically connects users with items, typically found on online platforms [1, 2, 3, 4, 5], assisting users in sorting and filtering items of interest from a vast selection [6]. In addition to recommending individual items to users, online service platforms such as Steam [^1] https://store.steampowered.com/ and Epic [^2] https://store.epicgames.com/, the digital distribution platforms for video games, and NetEase Cloud Music [^3] https://music.163.com/ and Spotify [^4] https://open.spotify.com/, the online music platforms, will also recommend bundles consisting of a number of items that have certain internal relationships. The prevalence of bundled items makes bundle recommendation an important task, and according to Deng et al.’s analysis of purchase statistics, more than 65.00% of game Love is Justice revenue in mobile games comes from these discounted bundles [7]. Therefore, bundle recommendation is an important part of the recommendation system. Bundle recommendation aims to improve the accuracy of the recommendation of bundles by mining the hidden messages of the interaction between users, items, and bundles.
Compared to traditional recommendation, the difficulty of bundle recommendation is how to utilize item elements including item-user interaction and item-bundle interaction as auxiliary information for bundle recommendation. This brings two problems, how to introduce item elements with item-user interactions and item-bundle interactions as auxiliary information, and how to utilize auxiliary information to improve bundle recommendation performance. While there has been a lot of work in the area of bundle recommendations [8, 9, 10, 11, 12, 13, 14, 15], scholars’ understanding of constructing a triadic relationship of user-item-bundle has been biased. The left part of the figure 1 shows the understanding of the previous work[8, 9, 12, 13], specifically, they simplified the ternary relationship by arguing that bundle recommendations can be generalized along two paths from users to bundles and users to items to bundles. This understanding is based on the user’s perspective, but it tears down the user-item-bundle ternary relationship. This way of understanding lengthens the information transmission path and brings information loss. We believe that users, items, and bundles are a kind of mutual connection relationship, and there is intra-connection within users and bundles, and the right part of the figure 1 illustrates our understanding. However, current research lacks a user-item-bundle triad construction that matches our understanding, and this construction also increases the complexity of the model.
Figure 1: Different understandings of the construction of the user-item-bundle triad: the right side is the understanding from previous work, and the left side is our understanding; where the blue squares represent users, the red squares represent items, and the green squares represent bundles; the grey arrows and the gold arrows represent the two paths from the understanding from previous work, and the black arrows and the bi-directional arrows represent the mutual connection; the blue curve and the green curve represent the intra connection of users and the intra connection of bundles, respectively.
To this end, we design the hypergraph-enhanced dual convolutional neural network (HED) model to address the problem of how to introduce item elements and how to utilise them. The model comprises three parts: embedding initialization layer, dual convolutional layer, and prediction layer. Based on the work [14], we constructed the hypergraph based on the mutual connections between users, projects and bundles, as well as the internal connections between users and bundles. We termed the hypergraph created as a complete hypergraph. The embedding initialization layer generates embedding vectors for each user, item, and bundle. The dual convolutional layer is capable of extracting information from the complete hypergraph and merging the information from the U-B graph produced by U-B interaction to improve the embedding vector and prediction. The layer is suggested based on the ultimate embedding vectors of both users and bundles. Experiments on NetEase and Youshu show that our proposed method is ahead of the state-of-the-art methods.
Overall, the main contributions of this paper are summarized as follows:
  • We construct a complete hypergraph to express the ternary relationship between users, items and bundles. The ternary relationship includes not only the mutual connections between users, items and bundles, but also the intra-connections within users and bundles.
  • We propose the HED model, which can mine the information contained in the complete hypergraph and utilize the information of the U-B graph to enhance the representation of users and bundles to achieve more accurate recommendations.
  • We conducted extensive experiments on two real-world datasets, and the results show that our proposed model outperforms the existing state-of-the-art baselines in terms of Recall and NDCG evaluation metrics in all three scenarios: Top 20, Top 40, and Top 80.
II Related Works
As an increasing number of scholars focus on bundle recommendation, the field has experienced rapid growth, giving rise to numerous advanced works. A portion of the work focuses on how to decouple the complex relationship between users, items, and bundles. For example, Chen et al. [16] aggregated item embeddings within a bundle using an attention mechanism to obtain a representation of the bundle, and jointly modeled user-bundle interactions and user-item interactions with a multitasking approach. Tan et al. [9] used a hierarchical approach to disentangle user preferences for items and bundle preferences. In contrast, Avny et al. [11] improved the Transformer structure to represent the data of users, items, and bundles, and achieved modeling of users’ preferences for each item in the bundle and for the whole bundle through a self-attention mechanism.
Influenced by graph neural networks in the field of recommender systems, many works have chosen to represent the relationships among users, items, and bundles with graph structures, transforming bundle recommendation into a link prediction problem on graphs. For example, Chang et al. [8] used two heterogeneous graphs to represent the three kinds of interactions among users, items, and bundles, and employed graph convolutional propagation to learn the user and bundle representations. Wang et al. [10] augmented the representations characterizing users and bundles with neighbor interactions. Ma et al. [12] represented the relationship between the three with bundle views and project views, and chose to use comparative learning to obtain user and bundle representation. Zhao et al. [13] learned the user and bundle representations by comparing the user intention in the global view and the user intention in the local view with contrast learning. Yu et al. [14] constructed the three kinds of interactions into a unified hypergraph and used a graph convolutional neural network to learn the user and bundle representations. Wei et al. [15] used the graph transformer to model and learn bundle strategies, and combined user preferences for items to obtain representations of users and bundles. These works aim to capture the three interactions involving users, items, and bundles using graph-based methodologies. In particular, the study by Yu et al. [14] utilized hypergraphs to represent these interactions. However, it has been observed that this representation is not comprehensive and does not sufficiently emphasize the role of items in bundle recommendations. To address this limitation, we have undertaken further refinement and construction of complete hypergraphs, building upon the foundation laid out in [14]. Our objective is to create a more detailed and unified structure that effectively encapsulates the interactions between users, items, and bundles, as well as the relationships within these three components.
III Problem Formulation
Consider a dataset with three main entities: bundles represented by the set $S_{B}=\{b_{i}\mid i\in 1,2,\dots,B\}$, users represented by the set $S_{U}=\{u_{j}\mid j\in 1,2,…,U\}$, and items represented by the set $S_{I}=\{i_{k}\mid k\in 1,2,\dots,I\}$. The interactions within this dataset are captured through three matrices:
  • User’s purchase interactions with bundles are represented by the matrix $A_{ub}$, where each entry $A_{ub}(i,j)$ signifies the user $u_{j}$’s interaction with bundle $b_{i}$.
  • User’s purchase interactions with individual items are captured by the matrix $A_{ui}$, with each entry $A_{ui}(j,k)$ indicating whether user $u_{j}$ has purchased item $i_{k}$.
  • The relationships between bundles and items are represented by the matrix $A_{bi}$, where each entry $A_{bi}(i,k)$ indicates whether bundle $b_{i}$ includes item $i_{k}$ as a subordinate component.
Given the assumptions, the task of bundle recommendation can be formulated as predicting the likelihood of purchasing bundles based on purchase interactions and subordinate interactions. Specifically, the model takes as inputs the matrices representing purchase interactions ($A_{ub}$ and $A_{ui}$), and subordinate interactions ($A_{bi}$). The output of the model is a set of personalized score values that indicate a user’s preference for purchasing bundles. The personalized scoring function is expressed as follows:
$\hat{y}=f(u,b~{}|~{}A_{ub},A_{ui},A_{bi};\theta)$ (1)
Here, $\theta$ represents the parameters of the model, and $\hat{y}$ is the personalized score function describing the likelihood of user $u$ purchasing bundle $b$. This score is used to portray the probability or strength of a user’s inclination to purchase a specific bundle. This formulation provides a consistent and clear representation of the bundle recommendation task and the associated scoring function.
IV Methodology
The formulated problem is addressed through the framework of the hypergraph-enhanced dual convolutional neural network (HED for short). The framework’s structure is depicted in Figure 2, highlighting key components such as the Embedding Initialization Layer, the Dual Convolutional Layer, and the Prediction Layer.
To elucidate the construction of the hypergraph in this paper, refer to Section IV-A. This section outlines the methodology for building the complete hypergraph that is central to the proposed model. Additionally, Section IV-B provides insights into the proposed graph neural network model.
Figure 2: The framework graph of HED: 1) Embedding initialization layer, which corresponds to subsection IV-B1 of the article; 2) Dual Convolutional Layer, which corresponds to subsection IV-B2 of the article; and 3) Prediction Layer, which corresponds to subsection IV-B3 of the article, blue squares represent the user, red squares represent the item, and green squares represent the bundle; rectangles of different colors represent the embedding vectors of different elements, solid arrows of different colors represent the data transfer of different elements, and solid arrows of different colors represent the data updating of different elements, blue represents user, red represents item, and green represents bundle, the construction method of the hypergraph $H$ corresponds to subsection IV-A of the article, the U-B graph is denoted as $h_{hb}$ in the article, the dashed box and ellipsis in the middle of the two images represent the omitted elements and interactions, the ellipsis in the embedding section represents the embedding corresponding to the omitted elements, and the dashed box in the Prediction Layer section represents the generated recommendation sequence..
IV-A Complete Hypergraph Structure
In the realm of recommendation systems, conventional simple graphs adeptly capture interactions between users and items through edges. However, in the context of bundle recommendation, the intricacies of representing users, items, and bundles necessitate an extension beyond the constraints of simple graphs. While simple graphs suffice for user-item interactions, they fall short in comprehensively modeling the relationships within the bundle recommendation domain.
To address this limitation, the work [14] introduced the concept of hypergraphs to bundle recommendation, offering a more expressive framework capable of connecting any number of vertices with a single hyperedge. The proposed unified hypergraph construction method in [14] lays a foundation for capturing the interplay between users, items, and bundles. However, it has notable gaps. Specifically, the unified hypergraph construction method overlooks the inherent interactions of each user and each bundle with itself. Additionally, it underplays crucial aspects such as the purchase interactions of users with items and the affiliation interactions of bundles with items.
Building upon the foundation laid by [14], this paper seeks to enhance hypergraph construction by introducing complete hypergraphs. Figure 3 provides a visual representation of these complete hypergraphs. In contrast to the unified hypergraph, our approach meticulously incorporates self-interactions of users and bundles, while also accounting for the intricate dynamics of user-item purchase interactions and bundle-item affiliation interactions.
Figure 3: Schematic representation of the hypergraph synthesis method: blue squares indicate users, red squares indicate items and green squares indicate bundles. Black arrows indicate computation and gold arrows indicate synthesis. By interacting $A_{ub}$, $A_{ui}$ and $A_{bb}$, U-B graph $h_{ub}$, U-I graph $h_{ui}$ and B-I graph $h_{bi}$ can be generated. The U-U graph (denoted as $h_{uu}$) and the B-B graph (denoted as $h_{bb}$) are calculated based on the interaction $h_{ub}$. In addition, $h_{ii}$ is the zero matrix and the ellipses represent omitted users, items, bundles, and corresponding interactions. The hypergraph $H$ is formed by the combination of these interaction graphs, denoted by the formula 2.
In the figure, the B-U graph also referred to as the B-U interaction matrix ($h_{ub}$), is derived from the interaction matrix $A_{ub}$. To clarify, the nodes in $h_{bu}$ correspond to the interactions between users and bundles. If $(i,j)$ is present in $A_{ub}$, it means that user $u_{i}$ purchased bundle $b_{j}$, so $h_{ub}(i,j)=1$, and vice versa $h_{ub}(i,j)=0$. Essentially, both $A_{ub}$ and $h_{ub}$ are mathematical representations of the relationship between a user and a bundle that generates a purchase, where $A_{ub}$ is represented by a pair of numbers representing the user number and the bundle number, and $h_{ub}$ is represented by the graph structure $h_{ub}(i,j)=1$. Similarly, the interaction $A_{ui}$ and $A_{bi}$ can be obtained as the U-I graph $h_{ui}$ and B-I graph $h_{bi}$. The matrix $h_{bu}$ is the transpose of the $h_{ub}$ matrix. Similarly, $h_{iu}$ and $h_{ib}$ are the transpose matrices of $h_{ui}$ and $h_{bi}$, respectively. The matrices $h_{uu}$ and $h_{bb}$ signify user-user and bundle-bundle interactions, respectively. These can be derived from the $h_{ub}$ interaction matrix. Specifically, for users A and B, if the number of bundles with the same interaction is greater than $n$, then users A and B are considered to have an interaction. This corresponds to $h_{uu}(a,b)$ and $h_{uu}(b,a)$ in the $h_{uu}$ matrix being set to 1, and vice versa. The parameter $n$ is experimentally tested under different datasets and is consistent with  [14]. Furthermore, it is important to note that each user is considered to interact with themselves. Consequently, we set the elements on the main diagonal of $h_{uu}$ to 1. The same principle applies to $h_{bb}$. Regarding $h_{bb}$, despite experimenting with construction methods similar to those used for $h_{uu}$ and attempting to set it as a unit matrix, the optimal experimental results were achieved when constructing $h_{bb}$ as a zero matrix. The complete hypergraph construction method can be represented by the formula 2, while the formula 3 is the construction method of the unified hypergraph [14]. In contrast to the unified hypergraph $H_{Unify}$, our construction of the full hypergraph $H$ emphasizes the interaction of items with bundles and users and can fully express the understanding of the right-hand side of the graph 1 concerning bundle recommendations.
$H=\begin{bmatrix}h_{uu}&h_{iu}&h_{bu}\\ h_{ui}&h_{ii}&h_{bi}\\ h_{ub}&h_{ib}&h_{bb}\\ \end{bmatrix}$ (2)
$H_{Unify}=\begin{bmatrix}h_{ui}&h_{uu}&h_{ub}\\ h_{bi}&h_{bu}&h_{bb}\\ \end{bmatrix}$ (3)
IV-B Network Structure
IV-B1 Embedding Initialization Layer
The focus of this paper lies in the domain of implicit feedback, where the available information is confined to the easily accessible history of interactions. Given this constraint, our model relies on the random initialization of embedding vectors for users, items, and bundles. These initialized embeddings constitute a crucial component of the trainable parameters within the model.
Building on the construction method outlined for the complete hypergraph, we arrange the embedding vectors to serve as inputs for the subsequent modules of the model. The stacking of these embedding vectors is carried out in accordance with the construction principles of the complete hypergraph.
Assuming that the embedding size of the model is set to $d$, the initialization generates an embedding of $E_{u}^{(1)}\in R^{U\times d}$ for users, $E_{i}^{(1)}\in R^{I\times d}$ for items, and $E_{b}^{(1)}\in R^{B\times d}$ for bundles, after the stacked embedding vector $E^{(1)}\in R^{(U+I+B)\times d}$, the formula for the stacking method is expressed as follows.
$E^{(1)}=\begin{bmatrix}E_{u}^{(1)}\\ E_{i}^{(1)}\\ E_{b}^{(1)}\\ \end{bmatrix}$ (4)
IV-B2 Dual Convolutional Layer
In a graph, messages are passed from nodes to edges and then from edges back to nodes[17][18]. Message passing is a paradigm for graph neural networks, specifically the propagation of features from one node to neighboring nodes. Similarly, in hypergraphs, messages are typically passed from nodes to hyperedges and then back to nodes. Although Feng et al. [19] proposed a hypergraph convolutional neural network, this network structure is not specific to recommender systems and does not take into account the characteristics of recommender systems. He et al. [20] demonstrated that, in recommender systems, feature transformations and nonlinear activations affect the model’s performance. Based on the previous work, the Efficient hypergraph convolution was designed in work [14].
$\begin{split}&\hat{H}=D_{V}^{-1/2}HD_{E}^{-1/2}H^{\text{T}}D_{V}^{-1/2}\\ &E^{(l+1)}=\hat{H}E^{(l)}\end{split}$ (5)
where $H$ is the hypergraph, $D_{V}$, and $D_{E}$ are the node degree matrices and edge degree matrices of the hypergraph $H$, respectively, where $\hat{H}$ refers to the Laplacian spectral normalization of hypergraph $H$, $E^{(l)}$ is the embedding matrix of layer $l$, and $E^{(l+1)}$ is the embedding matrix of layer $l+1$.
Although this design simplifies the hypergraph convolution operation and lowers the training cost, it neglects the features of bundle recommendation. The complete hypergraph introduces item elements, item-user interactions, and item-bundle interactions. Meanwhile, graph convolution operations on the complete hypergraph exploit auxiliary information but do not possess the module to return the auxiliary information for the task of bundle recommendation. We included $E_{i}$ to improve the learning of $E_{u}$ and $E_{b}$, yet the personalized scoring Equation 8 for the model does not feature $E_{i}$. As a result, we require new modules to enhance the expression of both $E_{u}$ and $E_{b}$, so that the features regarding an item in $E_{u}$ and $E_{b}$ will ultimately predict user-bundle interactions.
Therefore, we design the following convolution operation: the $l$-th layer embedding vector $E^{(l)}$ is computed as ${\bar{E}}^{(l)}$ after graph convolution on the complete hypergraph. We extract the embedding vectors of users in ${\bar{E}}^{(l)}$, ${\bar{E}}^{(l)}_{u}$ and bundles embedding vectors ${\bar{E}}^{(l)}_{b}$. ${\bar{E}}^{(l)}_{u}$ and ${\bar{E}}^{(l)}_{b}$ are respectively performed again on U-B graphs $h_{ub}$. The graph convolution is computed to obtain ${\tilde{E}}^{(l)}_{u}$ and ${\tilde{E}}^{(l)}_{b}$, while the output embedding $E^{(l+1)}$ is obtained by summing ${\bar{E}}^{(l)}$ with ${\tilde{E}}^{(l)}_{u}$ and ${\tilde{E}}^{(l)}_{b}$ according to parameters $\alpha$ and $\beta$. Since the item’s embedding does not participate in the convolution computation on the U-B graph, ${\bar{E}}^{(l)}_{i}$ is equal to $E_{i}^{(l+1)}$. Assuming that the complete hypergraph constructed is $H$, the graph consisting of the purchase interactions $h_{ub}$ is $h$. Then the convolution operation on the bipartite graph that we designed can be expressed by Equation 6.
$\begin{split}&\hat{H}=D_{V}^{-1/2}HD_{E}^{-1/2}H^{\text{T}}D_{V}^{-1/2}\\ &\hat{h}=D_{v}^{-1/2}hD_{e}^{-1/2}h^{\text{T}}D_{v}^{-1/2}\\ &{\bar{E}}^{(l)}=\hat{H}E^{(l)}=\begin{bmatrix}{\bar{E}}_{u}^{(l)}\\ {\bar{E}}_{i}^{(l)}\\ {\bar{E}}_{b}^{(l)}\\ \end{bmatrix}\\ &{\tilde{E}}^{(l)}_{u}=\hat{h}{\bar{E}}_{u}^{(l)}\\ &{\tilde{E}}^{(l)}_{b}=\hat{h}{\bar{E}}_{b}^{(l)}\\ &E^{(l+1)}=\begin{bmatrix}E_{u}^{(l+1)}\\ E_{i}^{(l+1)}\\ E_{b}^{(l+1)}\\ \end{bmatrix}=\begin{bmatrix}\alpha*{\bar{E}}_{u}^{(l)}+\beta*{\tilde{E}}^{(l)}_{u}\\ {\bar{E}}_{i}^{(l)}\\ \alpha*{\bar{E}}_{b}^{(l)}+\beta*{\tilde{E}}^{(l)}_{b}\\ \end{bmatrix}\end{split}$ (6)
Similarly, $D_{V}$, $D_{E},$ and $\hat{H}$ are the node degree matrix and edge degree matrix of the hypergraph $H$, Laplacian spectral normalization matrix, and $D_{v}$, $D_{e}$, and $\hat{h}$ are the node degree matrix and edge degree matrix of the U-B graph $h_{ub}$, Laplace spectral normalization matrix, respectively.
Since the model we designed contains multiple layers, only considering the embedding of the last layer will neglect the message of the neighbors mined by other layers. However, directly connecting the embedding of different layers will increase the model’s complexity. To address this problem, we draw inspiration from the method proposed in LightGCN [20], which combines the embedding of different layers by decaying them. This approach strategically integrates the embeddings from various layers, enabling the preservation and utilization of information mined at each level. Crucially, it achieves this without increasing the dimensionality of the embedding vectors, thereby maintaining the computational efficiency. And the combination method can be expressed by the formula, which is the final representation of users and bundles.
$\begin{split}E_{u}=\sum_{l=1}^{L}1/(l+1)*E_{u}^{(l)}\\ E_{b}=\sum_{l=1}^{L}1/(l+1)*E_{b}^{(l)}\end{split}$ (7)
where $L$ is the total number of layers.
IV-B3 Prediction Layer
After obtaining the final representation of users and bundles, the value of user $u$ personalized rating of bundle $b$ can be expressed as follows:
$\hat{y}_{ub}=e_{u}^{\text{T}}e_{b}$ (8)
where T represents the matrix transportation. The sequence of recommendations for bundling recommendations to user $u$ can be obtained by sorting the recommendations according to the personalized score values.
IV-B4 Model Training
The model leverages the User Interest Boundary (UIB) loss function, as proposed in the work by Zhuo et al. [21] to optimize its learning process. This loss function is specially designed to capture and reflect the nuances of users’ interests, making it particularly adept at handling the complexities inherent in user preferences modeling. The decision boundary established by UIB is formulated as follows:
$b_{u}=W^{\text{T}}E_{u}$ (9)
where $E_{u}$in $R^{d}$ denotes the embedding vector of user $u$ and $W$in $R_{d}$ is the learnable parameter. The UIB loss is a mixture of point-by-point loss and pairwise loss, which can be formulated as follows:
$\mathcal{L}=\sum_{(u,p)\in R^{+}}\phi(b_{u}-\hat{y}_{ub}^{p})+\sum_{(u,p)\in R^{-}}\phi(\hat{y}_{ub}^{n}-b_{u})$ (10)
Where $R^{+}$ represents the observed interactions set, $R^{-}$ represents the unobserved interactions set, $\hat{y}_{ub}^{p}$ represents the predicted scores of the observed interactions, and $\hat{y}_{ub}^{n}$ represents the unobserved interactions. Then, we use the AdamW optimizer [22] to minimize the loss function, which can iteratively update the weights of the neural network based on the training data.
V EXPERIMENTS
In this section, we delineate the comprehensive evaluation setup for our proposed approach, which we have named ‘HED’. The evaluation is meticulously designed to rigorously assess the efficacy, efficiency, and overall performance of ‘HED’ in various scenarios and against multiple benchmarks.
V-A Experiments Setup
Datasets. We choose the two most common datasets in bundle recommendation from different domains in the real world, Netease [^5] https://music.163.com/ and Youshu [^6] https://www.yousuu.com/, to validate our proposed model. Youshu is a book website from China, and the dataset provides interaction data of users, books, and bundles containing a series of books. The Netease dataset is constructed by crawling Netease Cloud Music, and it provides interaction data of users, music, and bundles containing a series of music. The parameters of the two datasets are shown in Table I.
TABLE I: Dataset statistics
  Dataset NetEase Youshu
  #U 18528 8039
#I 123628 32770
#B 22864 4771
#U-B 302303 51337
#U-I 1128065 138515
#B-I 1778838 176667
U-B Density 0.07% 0.13%
U-I Density 0.05% 0.05%
B-I Density 0.06% 0.11%
Evaluation Metrics. In evaluating the recommendation sequences generated by our model, we employ two of the most commonly used metrics in recommender systems: Recall at $k$ ($Recall@k$) and Normalized Discounted Cumulative Gain at $k$ ($NDCG@k$). These metrics are pivotal in gauging the effectiveness and accuracy of our recommendations. To ensure a comprehensive evaluation, we set the value of $k$ at three different levels: 20, 40, and 80.
Baselines. In the subsequent phase of our experiments, we conduct a comprehensive comparison between ‘HED’ and a selection of representative baseline methods. This comparative analysis is essential to validate the effectiveness and superior performance of ‘HED’. To ensure a fair and thorough comparison, we have chosen baseline methods that are widely recognized and have demonstrated robust performance in similar contexts. These methods encompass a range of techniques and algorithms, providing a diverse benchmark for assessing the strengths and potential areas of improvement for ‘HED’.
Next, we present these baseline methods:
  • DAM[16] is a neural network model utilizing attention mechanisms and multi-task learning techniques for bundled recommendations.
  • RGNN[10] utilizes graph neural network learning to construct user-bundle-item interaction graphs and bundle-item affiliation graphs, and introduces a multitasking learning framework to capture user’s preferences at the item level to augment recommendations for bundles.
  • MIDGN[13] is a graph neural network model capable of sorting out the combination of user intent and inter-bundle items at the global level, as well as separating the user’s intent from the items in each bundle at the local level.
  • CrossCBR[12] is a method of learning to connect the bundle view (U-B Graph) and the item view (U-I Graph & B-I Graph) through cross-view comparison. In addition, CrossCBR also increases the dispersion of different users/bundles, and the self-discrimination of the representation is increased.
  • UHBR[14] is a graph neural network model that unifies multiple associations between users, bundles, and items into a hypergraph and learns using graph convolutional neural networks.
  • BundleGT[15] is a graph transformer model that contains a token embedding layer, a hierarchical graph transformer layer, and a prediction layer
Training Details. In our experimental framework, we allocated 80% of the collected data for training ‘HED’, while the remaining 20% served as the validation set to assess the model’s performance. All experiments were carried out on a uniform hardware setup, featuring an Intel(R) Xeon(R) Platinum 8336C CPU and an Nvidia GeForce RTX 4090 GPU, ensuring consistency in computational performance across tests. Given the distinct characteristics of our datasets, we tailored the experimental parameters to optimize performance for each. Details are listed in Table II.
TABLE II: The parameter details.
NetEase Youshu
Embedding Size 128 64
$n$ 3 10
Epochs 500 300
Batch Size 2048 1024
$L$ 2
$\alpha$ 1/2
$\beta$ 1/100
Learning Rate 5e-3
Hypergraph Convolution Dropout 0.2
U-B graph Convolution Dropout 0.01
$L_{2}$-norm 0.1
V-B Comparative Results
In our comparative analysis, we endeavored to reproduce the baseline methods using publicly available source code or by closely adhering to the methodologies outlined in their respective publications. This approach was taken to ensure a fair and accurate comparison with our ‘HED’ model. However, it’s important to note that due to hardware constraints, the method proposed in Wei et al. (2023) [15] could not be effectively executed on the Netease dataset. Consequently, this limitation has resulted in the absence of data for Recall@80 and NDCG@80 metrics for this particular method. Table III presents a comprehensive performance comparison between ‘HED’ and the baseline methods.
TABLE III: Performance comparisons on two real-world datasets with the baselines.
Dataset Modle Recall@20 NDCG@20 Recall@40 NDCG@40 Recall@80 NDCG@80
NetEase DAM 0.0411 0.021 0.069 0.0281 0.109 0.0372
RGNN 0.0644 0.0327 0.1023 0.0431 0.1602 0.0551
MIDGN 0.0678 0.0343 0.1085 0.0451 0.1654 0.0578
CrossCBR 0.0842 0.0457 0.1264 0.0569 0.1805 0.0679
UHBR 0.0853 0.0452 0.1335 0.0577 0.1992 0.0732
BundleGT 0.0903 0.0478 0.1389 0.0606
HED 0.0963 0.051 0.1448 0.0635 0.212 0.0782
%Improv. 6.64% 6.69% 4.25% 4.79% 6.04% 6.39%
Youshu DAM 0.2082 0.1198 0.289 0.1418 0.3915 0.1658
RGNN 0.2732 0.1571 0.3814 0.1857 0.4937 0.2116
MIDGN 0.2682 0.1527 0.3712 0.1808 0.4817 0.2063
CrossCBR 0.2813 0.1668 0.3785 0.1938 0.4943 0.2198
UHBR 0.3049 0.1821 0.4166 0.2122 0.5308 0.2383
BundleGT 0.2905 0.1737 0.3909 0.2011 0.5124 0.2242
HED 0.3153 0.1903 0.423 0.2197 0.5377 0.2452
%Improv. 3.41% 4.50% 1.54% 3.53% 1.30% 2.90%
HED has become the new state-of-the-art model for bundled recommendation. According to table III, our proposed method is a generalized approach that outperforms the baseline methods of Youshu and NetEase by an average of 2.86% and 5.8%. We attribute this to the complete hypergraph’s superior ability to represent the ternary relationship between users, items, and bundles, and the dual convolution module’s capacity to extract information embedded in this relationship. This module also enhances the representation of users and bundles by leveraging information from the U-B graph, which significantly improves bundle recommendation task performance.
Through the joint analysis of the results of two datasets, we find that the improvements on NetEase are better than those on Youshu. The disparity in dataset size and interaction density between the Youshu and NetEase datasets could account for the observed differences. Specifically, the Youshu dataset has fewer users, items, and bundles than the NetEase dataset, as well as denser interactions among them, as indicated in Table I. This leads to the model’s $Recall@k$ and $NDCG@k$ metrics showing a significant improvement on the Youshu dataset compared to the NetEase dataset, resulting in a relatively smaller increase in performance on the former. However, in real-life datasets, there is often a similarity to the NetEase dataset where the number of users, items, and bundles is vast, and the interaction data between them is sparse. The proposed HED is applicable to practical situations.
The dual convolution module enhances the precision of the package proposition. Specifically, although HED performs better than previous models in both $Recall@k$ and $NDCG@k$, the improvement of $NDCG@k$ for the recommendation accuracy measure is greater than that of $Recall@k$, the maximum difference of 1.99%, the minimum difference of 0.05%. This may be due to the fact that previous work does not revert to the recommendation task after extracting information regarding the item. In contrast, our dual convolution module, which is designed to mine complete hypergraph information, can enhance the representation of users and bundles based on U-B graph information, thus improving the recommendation accuracy through the final personalized bundle recommendation score Formula 8.
V-C Ablation Study
In order to better explore the impact of the individual modules of the HED model on the whole, we disassemble it for Ablation Study. We implement four simplified variants of HED:
  • HED--c, removing the U-B graph convolution operation.
  • HED--cu, removing the U-B graph convolution and making $h_{uu}$ as a zero matrix.
  • HED--cb, removing the U-B graph convolution and making $h_{bb}$ as a zero matrix.
  • HED--cbu, removing the U-B graph convolution and making both $h_{uu}$ and $h_{bb}$ as zero matrices.
TABLE IV: Ablated models analysis on Youshu dataset.
Recall@20 NDCG@20 Recall@40 NDCG@40 Recall@80 NDCG@80
HED 0.316 (100%) 0.1892 (100%) 0.4209 (100%) 0.2175 (100%) 0.5395 (100%) 0.2445 (100%)
HED-c 0.3086 (97.66%) 0.1841 (97.3%) 0.4164 (98.93%) 0.2128 (97.84%) 0.5304 (98.31%) 0.2397 (98.04%)
HED-cu 0.2987 (94.53%) 0.1776 (93.87%) 0.4044 (96.08%) 0.2057 (94.57%) 0.5235 (97.03%) 0.2323 (95.01%)
HED-cb 0.2965 (93.83%) 0.1779 (94.03%) 0.4025 (95.63%) 0.2064 (94.90%) 0.5267 (97.63%) 0.2337 (95.58%)
HED-cbu 0.3011 (95.28%) 0.1792 (94.71%) 0.4046 (96.13%) 0.2076 (95.45%) 0.5254 (97.39%) 0.2347 (95.99%)
We summarize the results of our experiments in Table IV, where the leading number is the performance of the model, and the trailing percentage is the current performance as a percentage of the performance of the full model. With the table  IV we can get the following conclusions:
  • After excluding the dual convolution module, all six metrics of the model experience a varied decrease, where $NDCG@k$ metrics show a more significant decrease than $Recall@k$ metrics. This indicates that the dual convolution module can enhance the precision of bundle recommendation. Besides, the fact that the $Recall@20$ metric shows the most significant reduction among the three $Recall@k$ metrics confirms the same.
  • Removing the dual convolution module and setting $h_{uu}$ or $h_{bb}$ to the zero matrix leads to a significant drop in the overall model’s performance. Only two metrics, $Recall@20$, and $Recall@40$, show better results with $h_{uu}$ as the zero matrices, whereas the remaining four metrics approach better scores with $h_{bb}$ as the zero matrices. It is possible that the U-I interaction density is lower than the B-I interaction density, resulting in an under-represented complete hypergraph construction that impacts the final recommendation outcomes.
  • By removing the Dual Convolution Module and setting both $h_{uu}$ and $h_{bb}$ as zero matrices, the overall performance of the model deteriorates. However, it performs better than when only one of the matrices is zero. Specifically, all metrics are superior to the aforementioned cases, except for the $Recall@80$ metric, which is worse when $h_{bb}$ is a zero matrix. We analyzed the reason for this and concluded that creating both $h_{uu}$ and $h_{bb}$ zero matrices results in a more balanced hypergraph compared to the aforementioned cases. This construction places no emphasis on users or bundles, and hence, constructing $h_{uu}$ and $h_{bb}$ simultaneously is essential for improving the model’s performance. Conversely, constructing only one of them will result in an imbalanced hypergraph. Only one of them leads to a deviation from the complete hypergraph, thus diminishing the model’s performance.
V-D Sensitivity Analysis
In order to better explore the model’s sensitivity to each hyperparameter, we designed the Sensitivity Analysis experiment. Specifically, we experimented with scaling up the hyperparameters in the base experiment by a factor of $2$ or scaling down to $1/2$ of the original and summarized the results in Figures 4567.
Figure 4: Recall and NDCG performance under different $d$
Figure 5: Recall and NDCG performance under different $n$.
Figure 6: Recall and NDCG performance under different $\alpha$.
Figure 7: Recall and NDCG performance under different $\beta$
  • $d$ represents the dimension, which is the size of the embedding vectors for users, items, and bundles. The embedding vectors have a size of 64 for the baseline. The findings from the sensitivity analysis regarding the parameter $d$ are outlined in Figure 4, the left-hand line graph shows the recall metrics for different $d$ values, and the right-hand line graph shows the NDCG metrics for different $d$ values. When $d=16$, there is a more significant decline in model performance than the baseline. This is because, when $d$ is too small, the embedding vectors are insufficient in expressing the complete information acquired from the hypergraph and U-B graph. On the other hand, when $d>=32$, the embedding vectors exceed the dimensionality limit and can completely express the information derived from the complete hypergraph and U-B graph. The experiments support the claim that for $d>=32$, the model’s performance is similar to that of the baseline. Furthermore, the model performs better in terms of individual metrics. For instance, with $d=32$, all three metrics - $Recall@40$, $NDCG@40$, and $NDCG@80$ outperformed the baseline metrics.
  • The value of $n$ is crucial in constructing $h_{uu}$ and $h_{bb}$. For our baseline, $n$ is set at 10. The findings from the sensitivity analysis regarding the parameter $n$ are outlined in Figure 5, the left-hand line graph shows the recall metrics for different $n$ values, and the right-hand line graph shows the NDCG metrics for different $n$ values. If $n$ is doubled or halved, there is a slight reduction in model performance. Notably, the $NDCG@20$ metric experiences the highest decline of 1.48% at $n=5$. This suggests the impact of $n$ on model performance is perceptible but not significant. However, it is important to note that the optimal value of $n$ varies according to different datasets, and thus requires experimental determination.
  • $\alpha$ represents the combination coefficient of the complete hypergraph part embedding in the dual convolution module. The findings from the sensitivity analysis regarding the parameter $\alpha$ are outlined in Figure 6, the left-hand line graph shows the recall metrics for different $\alpha$ values, and the right-hand line graph shows the NDCG metrics for different $\alpha$ values. For our baseline, $\alpha$ is set at 1/2. The model’s performance experiences significant changes when $\alpha$ is increased to twice the original or decreased to half the original value. The significance of completing the convolution of the complete hypergraph lies in the valuable information of users, items, and bundles it provides for recommendation tasks. It must be noted that the complete hypergraph’s information is not adequately expressed when $\alpha=1/4$, whilst over-emphasized when $\alpha=1$. Therefore, it is advisable to set the value of $\alpha$ at an appropriate level that highlights the information present in the complete hypergraph as well as reflecting the information found in the U-B graph to some extent.
[…]
VI Conclusions
In this study, the significance of creating user-item-bundle triples in bundle recommendation is examined. A comprehensive hypergraph construction approach and HED model are introduced. A comprehensive hypergraph is a technique for conveying the ternary relationship formed by user, item, and bundle interactions. The HED model includes an embedding initialization layer, a dual convolutional layer, and a prediction layer. The embedding initialization layer produces embedding vectors for every user, item, and bundle. The dual convolutional layer extracts complete hypergraph information and enhances the embedding vectors by merging the U-B graph information. The prediction layer provides recommendations based on the final embedding vectors. A substantial volume of experiments conducted on Youshu and NetEase empirically prove the superiority of the HED model we have proposed.
Acknowledgment
This research is supported by the Fundamental Research Funds for the Central Universities (3102021ZDHQD09).

### v2

@@
-Hypergrah-Enhanced Dual Convolutional Network for Bundle Recommendation
+
 Abstract
-Bundle recommendations strive to offer users a set of items as a package named bundle, enhancing convenience and contributing to the seller’s revenue.
-While previous approaches have demonstrated notable performance, we argue that they may compromise the ternary relationship among users, items, and bundles.
+Bundle recommendations strive to offer users a set of items as a package named bundle, enhancing convenience and contributing to the distribution system’s revenue.
+Although existing bundle recommendation methodologies have demonstrated the ability to effectively capture user, item and bundle interactions, we posit that these approaches tend to compartmentalise the complex ternary relationship between users, items and bundles into two distinct paths, rather than considering it in its entirety.
 This compromise can result in information loss, ultimately impacting the overall model performance.
 To address this gap, we develop a unified model for bundle recommendation, termed Hypergraph-Enhanced Dual convolutional neural network (HED).
-Our approach is characterized by two key aspects.
-Firstly, we construct a complete hypergraph to capture interaction dynamics among users, items, and bundles.
-Secondly, we incorporate U-B interaction information to enhance the information representation derived from users and bundle embedding vectors.
+Three key aspects characterize our approach.
+First, we construct a complete hypergraph to capture the interactions dynamics between users, items, and bundles.
+Second, to enhance the quality of user and bundle embeddings and better utilize the complete hypergraph constructed, we generate embeddings for each item in addition to those for each user and bundle.
+Third, influenced by the coarse and fine ranking of recommender systems, we design a dual convolution module to achieve more accurate recommendations.
 Extensive experimental results on the Youshu and Netease datasets have demonstrated that HED surpasses state-of-the-art baselines, proving its effectiveness.
-In addition, various ablation studies and sensitivity analyses revealed the working mechanism and proved our effectiveness.
+Various ablation studies and sensitivity analyses revealed the working mechanism and proved our effectiveness.
+In addition, we discuss hypergraph construction and implementation efficiency.
 Codes and datasets are available at https://github.com/AAI-Lab/HED
-Index Terms:
-Bundle Recommendation, Graph Convolutional Network, Hypergraph, Recommender System
-I Introduction
-Recommender system is a type of system that automatically connects users with items, typically found on online platforms [1, 2, 3, 4, 5], assisting users in sorting and filtering items of interest from a vast selection [6].
-In addition to recommending individual items to users, online service platforms such as Steam [^1] https://store.steampowered.com/ and Epic [^2] https://store.epicgames.com/, the digital distribution platforms for video games, and NetEase Cloud Music [^3] https://music.163.com/ and Spotify [^4] https://open.spotify.com/, the online music platforms, will also recommend bundles consisting of a number of items that have certain internal relationships.
-The prevalence of bundled items makes bundle recommendation an important task, and according to Deng et al.’s analysis of purchase statistics, more than 65.00% of game Love is Justice revenue in mobile games comes from these discounted bundles [7].
-Therefore, bundle recommendation is an important part of the recommendation system.
-Bundle recommendation aims to improve the accuracy of the recommendation of bundles by mining the hidden messages of the interaction between users, items, and bundles.
-Compared to traditional recommendation, the difficulty of bundle recommendation is how to utilize item elements including item-user interaction and item-bundle interaction as auxiliary information for bundle recommendation.
-This brings two problems, how to introduce item elements with item-user interactions and item-bundle interactions as auxiliary information, and how to utilize auxiliary information to improve bundle recommendation performance.
-While there has been a lot of work in the area of bundle recommendations [8, 9, 10, 11, 12, 13, 14, 15], scholars’ understanding of constructing a triadic relationship of user-item-bundle has been biased.
-The left part of the figure 1 shows the understanding of the previous work[8, 9, 12, 13], specifically, they simplified the ternary relationship by arguing that bundle recommendations can be generalized along two paths from users to bundles and users to items to bundles.
-This understanding is based on the user’s perspective, but it tears down the user-item-bundle ternary relationship.
-This way of understanding lengthens the information transmission path and brings information loss.
-We believe that users, items, and bundles are a kind of mutual connection relationship, and there is intra-connection within users and bundles, and the right part of the figure 1 illustrates our understanding.
-However, current research lacks a user-item-bundle triad construction that matches our understanding, and this construction also increases the complexity of the model.
-Figure 1: Different understandings of the construction of the user-item-bundle triad: the right side is the understanding from previous work, and the left side is our understanding; where the blue squares represent users, the red squares represent items, and the green squares represent bundles; the grey arrows and the gold arrows represent the two paths from the understanding from previous work, and the black arrows and the bi-directional arrows represent the mutual connection; the blue curve and the green curve represent the intra connection of users and the intra connection of bundles, respectively.
-To this end, we design the hypergraph-enhanced dual convolutional neural network (HED) model to address the problem of how to introduce item elements and how to utilise them.
+keywords:
+[…]
+1 Introduction
+A recommender system is a crucial component of a Internet distribution system that rapidly searches for items, such as books, music, videos, and games, that may be of interest to users from a vast collection of items [21, 23, 22, 8, 52].
+The fundamental principle of recommender systems is to recommend items that are relevant to users, employing feature engineering techniques that utilise user preferences, item characteristics, and user interactions (purchases or clicks) [20].
+In addition to recommending individual items to users, online service platforms such as Steam [^1] https://store.steampowered.com/, Epic [^2] https://store.epicgames.com/, the digital distribution platforms for video games, and NetEase Cloud Music [^3] https://music.163.com/, Spotify [^4] https://open.spotify.com/, the online music platforms, will also recommend bundles consisting of a number of items that have certain internal relationships.
+The prevalence of bundles has made bundle recommendation an important task, according to Deng et al.’s analysis of purchase statistics, more than 65.00% of game Love is Justice revenue in mobile games comes from these discounted bundles [10].
+Bundle recommendation aims to improve the accuracy of the recommendation of bundles by mining the hidden messages of the interactions between users, items, and bundles.
+Compared with traditional recommendation, the difficulty of bundle recommendation is how to introduce and utilize item information to improve the performance of bundle recommendation.
+This brings up two questions, how to structure the user-item-bundle triadic relationship, and how to leverage the triadic relationship and focus on users and bundles for recommendations.
+While there has been a lot of work in the area of bundle recommendations [7, 32, 39, 2, 28, 51, 49, 42], scholars understanding of the triadic relationship of user-item-bundle has been biased.
+The left part of the figure 1 shows the understanding of the previous work[7, 32, 28, 51].
+They simplified the ternary relationship of bundles by arguing that bundle recommendations can be generalized along two paths: path 1 (users-bundles) and path 2 (users-items-bundles).
+This understanding artificially dismantles the user-item-bundle triadic relationship from the user’s point of view, lengthens the information transfer path and introduces information loss.
+We believe that users, items and bundles are in an interconnected relationship (both in terms of purchase and affiliation), and that there are internal connections within users or bundles, as illustrated by the right part of the figure 1.
+However, the current research does not conform to our understanding of the user-item-bundle ternary structure, nor does it have a bundle recommendation model based on this design.
+Figure 1: Different understandings of the construction of the user-item-bundle triadic relationship: the left side is the understanding from previous work, and the right side is our understanding.
+The blue, red, and green squares represent users, items, and bundles, respectively.
+The grey and gold arrows represent the two types of paths from the understanding from previous work, while the black and bi-directional arrows represent the mutual connection from our understanding.
+The blue and green curves represent the intra connection of users and the intra connection of bundles, respectively.
+To this end, we design the Hypergraph-Enhanced Dual convolutional neural network (HED) model.
 The model comprises three parts: Embedding initialization layer, Dual convolutional layer, and Prediction layer.
-Based on the work [14], we constructed the hypergraph based on the mutual connections between users, projects and bundles, as well as the internal connections between users and bundles.
-We termed the hypergraph created as a complete hypergraph.
-The embedding initialization layer generates embedding vectors for each user, item, and bundle.
-The dual convolutional layer is capable of extracting information from the complete hypergraph and merging the information from the U-B graph produced by U-B interaction to improve the embedding vector and prediction.
-The layer is suggested based on the ultimate embedding vectors of both users and bundles.
+Based on the work [49], we constructed hypergraphs called complete hypergraphs based on the mutual connections between users, items, and bundles as well as the internal connections between users or bundles.
+The Embedding initialization layer generates embeddings for each user, item, and bundle, with the aim of improving the quality of user and bundle embeddings through item embeddings.
+The Dual convolutional layer is composed of two parts.
+The first part convolves the hypergraph to obtain information about the user, item, and bundle from the complete hypergraph.
+The second part convolves the U-B graph, which is obtained from the U-B interactions, to obtain information about the user and the bundle.
+The final embedding of the user and the bundle is obtained by combining the information from both parts.
+The Prediction layer is suggested based on the ultimate embedding of both users and bundles.
 Experiments on NetEase and Youshu show that our proposed method is ahead of the state-of-the-art methods.
 Overall, the main contributions of this paper are summarized as follows:
-  • We construct a complete hypergraph to express the ternary relationship between users, items and bundles.
-The ternary relationship includes not only the mutual connections between users, items and bundles, but also the intra-connections within users and bundles.
+  • We construct a complete hypergraph to express the ternary relationship between users, items, and bundles.
+The ternary relationship includes not only the mutual connections between users, items, and bundles but also the intra connections within users or bundles.
   • We propose the HED model, which can mine the information contained in the complete hypergraph and utilize the information of the U-B graph to enhance the representation of users and bundles to achieve more accurate recommendations.
   • We conducted extensive experiments on two real-world datasets, and the results show that our proposed model outperforms the existing state-of-the-art baselines in terms of Recall and NDCG evaluation metrics in all three scenarios: Top 20, Top 40, and Top 80.
-II Related Works
+Furthermore, the increase in NDCG metrics and the metrics of Top 20 demonstrates that the HED model enhances the precision of recommendations.
+2 Related Works
+This section presents a review of related work in three areas: (1) Graph-Based Recommender System, (2) Hypergraph-Based Recommender System, and (3) Bundle Recommendation.
+2.1 Graph-Based Recommender System
+A recommender system is a system that learns from a user’s historical interactions with items and subsequently recommends items to the user that may be of interest to them.
+The user and the item can be regarded as nodes of a graph, with the interaction information between them forming the edges of the graph.
+This interaction information is thus well-suited to expression in the form of a graph.
+The development of graph neural networks has led to the clarification of several key paradigms in the field of graph analysis, including link prediction, graph classification, graph generation, and graph node classification.
+Additionally, the emergence of classical graph neural network algorithms, such as Spectral CNN[5], GCN[19], GraphSAGE[16], and GAT[33], has facilitated the development of graph neural network-based recommendation models.
+In the field of personalised recommendation, He et al.
+innovatively constructed the interaction information of users and items as a bipartite graph, and designed the Neural Graph Collaborative Filtering framework, which learns the higher-order information of the bipartite graph and updates it [37].
+The embedding of users and items through GCN.
+NGCF not only demonstrates outstanding performance, but also sets a precedent for GNN-based recommender systems.
+He et al.
+demonstrated that feature transformation and nonlinear activation can negatively affect GCNs through ablation analysis, which led to the development of the far-reaching LightGCN model as the research progressed [17].
+Liu et al.
+decomposed the embedding into multiple segments related to semantics, thereby enhancing the performance and interpretability of the model [25].
+Wu et al.
+enhanced accuracy and robustness through self-supervised learning based on LightGCN, thereby alleviating the long-tail problem [43].
+Xia et al.
+proposed an incremental graph convolutional network (GCN) as a means of addressing the cold-start problem inherent to GCN when applied to new users and new items [44].
+Lin et al.
+enhanced the efficacy of recommendation algorithms by integrating contrast learning, which categorizes nodes and their structural neighbors as positive contrast pairs [24].
+The prevalence of graph neural networks is not only limited to the field of personalised recommendation, but has also flourished in the fields of hybrid recommendation  [15], session recommendation  [31], point-of-interest (POIs) recommendation  [41], group recommendation  [1], bundle recommendation [28], and other segmented recommender system domains have also flourished.
+Gatzioura et al.
+developed a graph-based model that incorporates content-based and collaborative filtering-based models for mining music descriptions with user preferences for music recommendation [15].
+Sheng et al.
+constructed interactions and sessions as Weighted Global Item Graph and Local Session Graph, respectively, and designed an Enhanced Graph Neural Network to mine the information of the two graphs for session recommendation [31].
+Wang et al.
+put forward the Adaptive Graph Representation-enhanced Attention Network (AGRAN) framework with a view to learning spatio-temporal information of POI and the user’s preferences for POI recommendation, using a combination of techniques including graph structure learning and the construction of an adaptive POI graph matrix [41].
+Abolghasemi et al.
+evaluated the similarity of user graphs using a graph convolutional neural network and designed the GcPp technique to cluster highly similar users into clusters, which in turn enables group recommendation [1].
+Furthermore, information external to users and items can be represented as a knowledge graph, which in turn facilitates the generation of more accurate recommendations based on historical interactions [13].
+2.2 Hypergraph-Based Recommender System
+As the study progressed, the use of simple graphs was no longer sufficient to convey the required information.
+Hypergraph has attracted significant attention due to its capacity to accommodate diverse data types and establish intricate connections, offering a promising avenue for addressing the challenges posed by such vast and heterogeneous datasets [4].
+The proposal of hypergraph neural networks has led to the emergence of a growing body of research on hypergraph-based approaches in the field of recommender systems [12, 14].
+This includes studies on personalized recommendation [35, 18, 45, 40, 27], POI recommendation [47], session-based recommendation [46, 11], group recommendation [50, 36], and social recommendation [48].
+In a personalized recommendation task, Wang et al.
+constructed a series of hypergraphs based on item correlations for different time periods and employed a Hypergraph Convolution Network combined with self-attention to learn hypergraph information for recommendation [35].
+The majority of works [18, 45, 40, 27] construct hypergraphs based on users and items.
+Ji et al.
+designed a two-channel collaborative filtering framework to learn hypergraph information [18].
+Xia et al.
+constructed hypergraphs based on users and items and made recommendations through contrastive learning [45].
+Wang et al.
+constructed hypergraphs based on users and items and made recommendations through attention and dynamics clustering [40].
+Ma et al.
+constructed a hypergraph based on users and items and learned hypergraph information through cross-view hypergraph contrastive learning for recommendation [27].
+In the session-based recommendation task, scholars construct hypergraphs based on sessions and items, Xia et al.
+learn hypergraph information through a two-channel hypergraph convolutional network and train the model through a self-supervised task [46], and Ding et al.
+learn two item representations through Hypergraph Convolutional Networks (HGCN) and Gate Recursive Units (GRU) with an attentional mechanism for recommendation [11].
+In the social recommendation task, Yu et al.
+constructed a hypergraph based on social network and user-item interaction graph, and used self-supervised learning to train the hypergraph convolutional network [48].
+In the POI recommendation task, Yan et al.
+constructed a trajectory-based hypergraph and proposed a spatio-temporal hypergraph convolutional network that fuses complex higher-order information and global collaborative relationships among trajectories for recommendation [47].
+2.3 Bundle Recommendation
 As an increasing number of scholars focus on bundle recommendation, the field has experienced rapid growth, giving rise to numerous advanced works.
 A portion of the work focuses on how to decouple the complex relationship between users, items, and bundles.
-For example, Chen et al.
+For example, Rendle et al.
+treat bundles as atomic, use matrix decomposition to learn preferences, and make recommendations via Markov chains [30].
+Similarly, Cao et al.
+employed a factor decomposition model to capture user preferences for items and bundles and to make recommendations based on co-occurring information about items and bundles [6].
+In addition, the method they developed addresses the cold-start problem for new items.
+Chen et al.
 [9] aggregated item embeddings within a bundle using an attention mechanism to obtain a representation of the bundle, and jointly modeled user-bundle interactions and user-item interactions with a multitasking approach.
@@
 [2] improved the Transformer structure to represent the data of users, items, and bundles, and achieved modeling of users’ preferences for each item in the bundle and for the whole bundle through a self-attention mechanism.
+Figure 2: The framework graph of HED: 1) Embedding initialization layer, which corresponds to subsection 4.2.1 of the article; 2) Dual Convolutional Layer, which corresponds to subsection 4.2.2 of the article; and 3) Prediction Layer, which corresponds to subsection 4.2.3 of the article.
+The blue, red, and green squares represent users, items, and bundles, respectively.
+Rectangles of different colors represent the embedding of different elements.
+Data transmission and updating are shown by solid and dashed lines.
+The hypergraph H is the complete hypergraph constructed by subsection 4.1.
+The dashed boxes in the prediction layer section indicate the generated recommendation sequences.
 Influenced by graph neural networks in the field of recommender systems, many works have chosen to represent the relationships among users, items, and bundles with graph structures, transforming bundle recommendation into a link prediction problem on graphs.
@@
 Ma et al.
-[12] represented the relationship between the three with bundle views and project views, and chose to use comparative learning to obtain user and bundle representation.
+[28] represented the relationship between the three with bundle views and item views, and chose to use comparative learning to obtain user and bundle representation.
 Zhao et al.
@@
 [49] utilized hypergraphs to represent these interactions.
-However, it has been observed that this representation is not comprehensive and does not sufficiently emphasize the role of items in bundle recommendations.
-To address this limitation, we have undertaken further refinement and construction of complete hypergraphs, building upon the foundation laid out in [14].
-Our objective is to create a more detailed and unified structure that effectively encapsulates the interactions between users, items, and bundles, as well as the relationships within these three components.
-III Problem Formulation
-Consider a dataset with three main entities: bundles represented by the set $S_{B}=\{b_{i}\mid i\in 1,2,\dots,B\}$, users represented by the set $S_{U}=\{u_{j}\mid j\in 1,2,…,U\}$, and items represented by the set $S_{I}=\{i_{k}\mid k\in 1,2,\dots,I\}$.
-The interactions within this dataset are captured through three matrices:
-  • User’s purchase interactions with bundles are represented by the matrix $A_{ub}$, where each entry $A_{ub}(i,j)$ signifies the user $u_{j}$’s interaction with bundle $b_{i}$.
-  • User’s purchase interactions with individual items are captured by the matrix $A_{ui}$, with each entry $A_{ui}(j,k)$ indicating whether user $u_{j}$ has purchased item $i_{k}$.
-  • The relationships between bundles and items are represented by the matrix $A_{bi}$, where each entry $A_{bi}(i,k)$ indicates whether bundle $b_{i}$ includes item $i_{k}$ as a subordinate component.
-Given the assumptions, the task of bundle recommendation can be formulated as predicting the likelihood of purchasing bundles based on purchase interactions and subordinate interactions.
-Specifically, the model takes as inputs the matrices representing purchase interactions ($A_{ub}$ and $A_{ui}$), and subordinate interactions ($A_{bi}$).
+However, this approach is not complete enough.
+Although items are introduced, the hypergraph underutilizes the interaction information for the included items resulting in an incompletely constructed hypergraph, and there is no generation of embedding for items to assist users and bundles in learning the hypergraph structure, as well as a lack of highlighting users interacting with bundles after learning the hypergraph.
+These factors not only led to poor model performance with the larger and more complex NetEase dataset but also led to the appearance of bundles that users wanted to interact with in a backward position.
+3 Problem Formulation
+Consider a dataset with three main entities: bundles represented by the set $S_{B}=\{b_{i}\mid i\in 1,2,…,B\}$, users represented by the set $S_{U}=\{u_{j}\mid j\in 1,2,…,U\}$, and items represented by the set $S_{I}=\{i_{k}\mid k\in 1,2,…,I\}$.
+The interactions within this dataset are captured through three matrices: User’s purchase interactions with bundles are represented by the matrix $A_{ub}$, where each entry $A_{ub}(i,j)$ signifies the user $u_{j}$’s interactions with bundle $b_{i}$.
+The same reasoning can be used to obtain the purchase interactions of users with individual items $A_{ui}$ and the affiliation interactions of bundles and items $A_{bi}$.
+Given the assumptions, the task of bundle recommendation can be formulated as predicting the likelihood of purchasing bundles based on purchase interactions and affiliation interactions.
+Specifically, the model takes as inputs the matrices representing purchase interactions ($A_{ub}$ and $A_{ui}$), and affiliation interactions ($A_{bi}$).
 The output of the model is a set of personalized score values that indicate a user’s preference for purchasing bundles.
 The personalized scoring function is expressed as follows:
-$\hat{y}=f(u,b~{}|~{}A_{ub},A_{ui},A_{bi};\theta)$ (1)
+Figure 3: Schematic diagram illustrating the structure of the hypergraph: blue squares indicate users, red squares indicate items and green squares indicate bundles.
+Black arrows indicate computation and gold arrows indicate synthesis.
+By interacting $A_{ub}$, $A_{ui}$ and $A_{bb}$, U-B graph $h_{ub}$, U-I graph $h_{ui}$ and B-I graph $h_{bi}$ can be generated.
+The U-U graph (denoted as $h_{uu}$) and the B-B graph (denoted as $h_{bb}$) are calculated based on the interactions $h_{ub}$.
+In addition, $h_{ii}$ is the zero matrix and the ellipses represent omitted users, items, bundles, and corresponding interactions.
+The hypergraph $H$ is formed by the combination of these interaction graphs, denoted by the Equation (2).
 Here, $\theta$ represents the parameters of the model, and $\hat{y}$ is the personalized score function describing the likelihood of user $u$ purchasing bundle $b$.
@@
 This formulation provides a consistent and clear representation of the bundle recommendation task and the associated scoring function.
-IV Methodology
-The formulated problem is addressed through the framework of the hypergraph-enhanced dual convolutional neural network (HED for short).
+4 Methodology
+The formulated problem is addressed through the framework of the HED.
 The framework’s structure is depicted in Figure 2, highlighting key components such as the Embedding Initialization Layer, the Dual Convolutional Layer, and the Prediction Layer.
-To elucidate the construction of the hypergraph in this paper, refer to Section IV-A.
+The construction of the hypergraph is explained in section 4.1.
 This section outlines the methodology for building the complete hypergraph that is central to the proposed model.
-Additionally, Section IV-B provides insights into the proposed graph neural network model.
-Figure 2: The framework graph of HED: 1) Embedding initialization layer, which corresponds to subsection IV-B1 of the article; 2) Dual Convolutional Layer, which corresponds to subsection IV-B2 of the article; and 3) Prediction Layer, which corresponds to subsection IV-B3 of the article, blue squares represent the user, red squares represent the item, and green squares represent the bundle; rectangles of different colors represent the embedding vectors of different elements, solid arrows of different colors represent the data transfer of different elements, and solid arrows of different colors represent the data updating of different elements, blue represents user, red represents item, and green represents bundle, the construction method of the hypergraph $H$ corresponds to subsection IV-A of the article, the U-B graph is denoted as $h_{hb}$ in the article, the dashed box and ellipsis in the middle of the two images represent the omitted elements and interactions, the ellipsis in the embedding section represents the embedding corresponding to the omitted elements, and the dashed box in the Prediction Layer section represents the generated recommendation sequence..
-IV-A Complete Hypergraph Structure
+Additionally, Section 4.2 provides insights into the proposed graph neural network model.
+4.1 Complete Hypergraph Structure
 In the realm of recommendation systems, conventional simple graphs adeptly capture interactions between users and items through edges.
@@
 While simple graphs suffice for user-item interactions, they fall short in comprehensively modeling the relationships within the bundle recommendation domain.
-To address this limitation, the work [14] introduced the concept of hypergraphs to bundle recommendation, offering a more expressive framework capable of connecting any number of vertices with a single hyperedge.
-The proposed unified hypergraph construction method in [14] lays a foundation for capturing the interplay between users, items, and bundles.
+$\hat{y}=f(u,b~{}|~{}A_{ub},A_{ui},A_{bi};\theta)$ (1)
+To address this limitation, the work [49] proposes a unified hypergraph construction methodology that introduces the concept of hypergraphs to bundle recommendations, laying the groundwork for capturing interactions between users, items, and bundles.
 However, it has notable gaps.
 Specifically, the unified hypergraph construction method overlooks the inherent interactions of each user and each bundle with itself.
-Additionally, it underplays crucial aspects such as the purchase interactions of users with items and the affiliation interactions of bundles with items.
-Building upon the foundation laid by [14], this paper seeks to enhance hypergraph construction by introducing complete hypergraphs.
+Additionally, it underplays crucial aspects such as the purchase interactions of users with items and the affiliation interactions of bundles with items.Building upon the foundation , this paper seeks to enhance hypergraph construction by introducing complete hypergraphs.
 Figure 3 provides a visual representation of these complete hypergraphs.
-In contrast to the unified hypergraph, our approach meticulously incorporates self-interactions of users and bundles, while also accounting for the intricate dynamics of user-item purchase interactions and bundle-item affiliation interactions.
-Figure 3: Schematic representation of the hypergraph synthesis method: blue squares indicate users, red squares indicate items and green squares indicate bundles.
-Black arrows indicate computation and gold arrows indicate synthesis.
-By interacting $A_{ub}$, $A_{ui}$ and $A_{bb}$, U-B graph $h_{ub}$, U-I graph $h_{ui}$ and B-I graph $h_{bi}$ can be generated.
-The U-U graph (denoted as $h_{uu}$) and the B-B graph (denoted as $h_{bb}$) are calculated based on the interaction $h_{ub}$.
-In addition, $h_{ii}$ is the zero matrix and the ellipses represent omitted users, items, bundles, and corresponding interactions.
-The hypergraph $H$ is formed by the combination of these interaction graphs, denoted by the formula 2.
+In contrast to the unified hypergraph, our approach meticulously incorporates self-interactions of users or bundles, while also accounting for the intricate dynamics of user-item purchase interactions and bundle-item affiliation interactions.
+$H=\begin{bmatrix}h_{uu}&h_{iu}&h_{bu}\\ h_{ui}&h_{ii}&h_{bi}\\ h_{ub}&h_{ib}&h_{bb}\\ \end{bmatrix}$ (2)
 In the figure, the B-U graph also referred to as the B-U interaction matrix ($h_{ub}$), is derived from the interaction matrix $A_{ub}$.
-To clarify, the nodes in $h_{bu}$ correspond to the interactions between users and bundles.
+To clarify, the nodes in $h_{bu}$ correspond to the interactions between users or bundles.
 If $(i,j)$ is present in $A_{ub}$, it means that user $u_{i}$ purchased bundle $b_{j}$, so $h_{ub}(i,j)=1$, and vice versa $h_{ub}(i,j)=0$.
-Essentially, both $A_{ub}$ and $h_{ub}$ are mathematical representations of the relationship between a user and a bundle that generates a purchase, where $A_{ub}$ is represented by a pair of numbers representing the user number and the bundle number, and $h_{ub}$ is represented by the graph structure $h_{ub}(i,j)=1$.
-Similarly, the interaction $A_{ui}$ and $A_{bi}$ can be obtained as the U-I graph $h_{ui}$ and B-I graph $h_{bi}$.
-The matrix $h_{bu}$ is the transpose of the $h_{ub}$ matrix.
-Similarly, $h_{iu}$ and $h_{ib}$ are the transpose matrices of $h_{ui}$ and $h_{bi}$, respectively.
 The matrices $h_{uu}$ and $h_{bb}$ signify user-user and bundle-bundle interactions, respectively.
 These can be derived from the $h_{ub}$ interaction matrix.
-Specifically, for users A and B, if the number of bundles with the same interaction is greater than $n$, then users A and B are considered to have an interaction.
+Specifically, for users A and B, if the number of bundles with the same mutual interaction is greater than $n$, then users A and B are considered to have an intra interaction.
 This corresponds to $h_{uu}(a,b)$ and $h_{uu}(b,a)$ in the $h_{uu}$ matrix being set to 1, and vice versa.
@@
 The same principle applies to $h_{bb}$.
-Regarding $h_{bb}$, despite experimenting with construction methods similar to those used for $h_{uu}$ and attempting to set it as a unit matrix, the optimal experimental results were achieved when constructing $h_{bb}$ as a zero matrix.
-The complete hypergraph construction method can be represented by the formula 2, while the formula 3 is the construction method of the unified hypergraph [14].
-In contrast to the unified hypergraph $H_{Unify}$, our construction of the full hypergraph $H$ emphasizes the interaction of items with bundles and users and can fully express the understanding of the right-hand side of the graph 1 concerning bundle recommendations.
-$H=\begin{bmatrix}h_{uu}&h_{iu}&h_{bu}\\ h_{ui}&h_{ii}&h_{bi}\\ h_{ub}&h_{ib}&h_{bb}\\ \end{bmatrix}$ (2)
-$H_{Unify}=\begin{bmatrix}h_{ui}&h_{uu}&h_{ub}\\ h_{bi}&h_{bu}&h_{bb}\\ \end{bmatrix}$ (3)
-IV-B Network Structure
-IV-B1 Embedding Initialization Layer
+Regarding $h_{ii}$, despite experimenting with construction methods similar to those used for $h_{uu}$ and attempting to set it as a unit matrix, the optimal experimental results were achieved when constructing $h_{ii}$ as a zero matrix.
+The specific analysis can be seen in 5.5.
+The complete hypergraph construction method can be represented by the Equation (2), our construction of the full hypergraph $H$ emphasizes the interactions of items with bundles and users and can fully express the understanding of the right-hand side of the graph 1 concerning bundle recommendations.
+4.2 Network Structure
+4.2.1 Embedding initialization layer
 The focus of this paper lies in the domain of implicit feedback, where the available information is confined to the easily accessible history of interactions.
-Given this constraint, our model relies on the random initialization of embedding vectors for users, items, and bundles.
+Given this constraint, our model relies on the random initialization of embedding for users, items, and bundles.
 These initialized embeddings constitute a crucial component of the trainable parameters within the model.
-Building on the construction method outlined for the complete hypergraph, we arrange the embedding vectors to serve as inputs for the subsequent modules of the model.
-The stacking of these embedding vectors is carried out in accordance with the construction principles of the complete hypergraph.
-Assuming that the embedding size of the model is set to $d$, the initialization generates an embedding of $E_{u}^{(1)}\in R^{U\times d}$ for users, $E_{i}^{(1)}\in R^{I\times d}$ for items, and $E_{b}^{(1)}\in R^{B\times d}$ for bundles, after the stacked embedding vector $E^{(1)}\in R^{(U+I+B)\times d}$, the formula for the stacking method is expressed as follows.
+Building on the construction method outlined for the complete hypergraph, we arrange the embedding to serve as inputs for the subsequent modules of the model.
+The stacking of these embedding is carried out in accordance with the construction principles of the complete hypergraph.
+Assuming that the embedding size of the model is set to $d$, the initialization generates an embedding of $E_{u}^{(1)}\in R^{U\times d}$ for users, $E_{i}^{(1)}\in R^{I\times d}$ for items, and $E_{b}^{(1)}\in R^{B\times d}$ for bundles, after the stacked embedding vector $E^{(1)}\in R^{(U+I+B)\times d}$, the equation for the stacking method is expressed as follows.
 $E^{(1)}=\begin{bmatrix}E_{u}^{(1)}\\ E_{i}^{(1)}\\ E_{b}^{(1)}\\ \end{bmatrix}$ (3)
-IV-B2 Dual Convolutional Layer
-In a graph, messages are passed from nodes to edges and then from edges back to nodes[17][18].
+4.2.2 Dual convolutional layer
+In a graph, messages are passed from nodes to edges and then from edges back to nodes[kipf2016semi][38].
 Message passing is a paradigm for graph neural networks, specifically the propagation of features from one node to neighboring nodes.
@@
 where $H$ is the hypergraph, $D_{V}$, and $D_{E}$ are the node degree matrices and edge degree matrices of the hypergraph $H$, respectively, where $\hat{H}$ refers to the Laplacian spectral normalization of hypergraph $H$, $E^{(l)}$ is the embedding matrix of layer $l$, and $E^{(l+1)}$ is the embedding matrix of layer $l+1$.
-Although this design simplifies the hypergraph convolution operation and lowers the training cost, it neglects the features of bundle recommendation.
+Although this design simplifies the hypergraph convolution operation and reduces the training cost, it ignores the characteristics of bundled recommendation.
 The complete hypergraph introduces item elements, item-user interactions, and item-bundle interactions.
-Meanwhile, graph convolution operations on the complete hypergraph exploit auxiliary information but do not possess the module to return the auxiliary information for the task of bundle recommendation.
-We included $E_{i}$ to improve the learning of $E_{u}$ and $E_{b}$, yet the personalized scoring Equation 8 for the model does not feature $E_{i}$.
-As a result, we require new modules to enhance the expression of both $E_{u}$ and $E_{b}$, so that the features regarding an item in $E_{u}$ and $E_{b}$ will ultimately predict user-bundle interactions.
+Meanwhile, graph convolution operations on complete hypergraphs use auxiliary information, but do not have modules that focus on bundle recommendations.
+We include $E_{i}$ to improve the learning of $E_{u}$ and $E_{b}$, but the model’s personalised scoring Equation (7) does not include $E_{i}$.
+As a result, we need new modules that utilise $E_{i}$ to enhance the expressiveness of $E_{u}$ and $E_{b}$ so that $E_{u}$ and $E_{b}$ can more accurately predict the user’s interactions with the bundle.
 Therefore, we design the following convolution operation: the $l$-th layer embedding vector $E^{(l)}$ is computed as ${\bar{E}}^{(l)}$ after graph convolution on the complete hypergraph.
-We extract the embedding vectors of users in ${\bar{E}}^{(l)}$, ${\bar{E}}^{(l)}_{u}$ and bundles embedding vectors ${\bar{E}}^{(l)}_{b}$.
+We extract the embedding of users in ${\bar{E}}^{(l)}$, ${\bar{E}}^{(l)}_{u}$ and bundles embedding ${\bar{E}}^{(l)}_{b}$.
 ${\bar{E}}^{(l)}_{u}$ and ${\bar{E}}^{(l)}_{b}$ are respectively performed again on U-B graphs $h_{ub}$.
@@
 Assuming that the complete hypergraph constructed is $H$, the graph consisting of the purchase interactions $h_{ub}$ is $h$.
-Then the convolution operation on the bipartite graph that we designed can be expressed by Equation 6.
+Then the convolution operation on the bipartite graph that we designed can be expressed by Equation (5).
 $\begin{split}&\hat{H}=D_{V}^{-1/2}HD_{E}^{-1/2}H^{\text{T}}D_{V}^{-1/2}\\ &\hat{h}=D_{v}^{-1/2}hD_{e}^{-1/2}h^{\text{T}}D_{v}^{-1/2}\\ &{\bar{E}}^{(l)}=\hat{H}E^{(l)}=\begin{bmatrix}{\bar{E}}_{u}^{(l)}\\ {\bar{E}}_{i}^{(l)}\\ {\bar{E}}_{b}^{(l)}\\ \end{bmatrix}\\ &{\tilde{E}}^{(l)}_{u}=\hat{h}{\bar{E}}_{u}^{(l)}\\ &{\tilde{E}}^{(l)}_{b}=\hat{h}{\bar{E}}_{b}^{(l)}\\ &E^{(l+1)}=\begin{bmatrix}E_{u}^{(l+1)}\\ E_{i}^{(l+1)}\\ E_{b}^{(l+1)}\\ \end{bmatrix}=\begin{bmatrix}\alpha*{\bar{E}}_{u}^{(l)}+\beta*{\tilde{E}}^{(l)}_{u}\\ {\bar{E}}_{i}^{(l)}\\ \alpha*{\bar{E}}_{b}^{(l)}+\beta*{\tilde{E}}^{(l)}_{b}\\ \end{bmatrix}\end{split}$ (5)
@@
 This approach strategically integrates the embeddings from various layers, enabling the preservation and utilization of information mined at each level.
-Crucially, it achieves this without increasing the dimensionality of the embedding vectors, thereby maintaining the computational efficiency.
-And the combination method can be expressed by the formula, which is the final representation of users and bundles.
+Crucially, it achieves this without increasing the dimensionality of the embedding, thereby maintaining the computational efficiency.
+The combination method can be expressed by the Equation (6), which is the final representation of users and bundles.
 $\begin{split}E_{u}=\sum_{l=1}^{L}1/(l+1)*E_{u}^{(l)}\\ E_{b}=\sum_{l=1}^{L}1/(l+1)*E_{b}^{(l)}\end{split}$ (6)
 where $L$ is the total number of layers.
-IV-B3 Prediction Layer
+4.2.3 Prediction layer
 After obtaining the final representation of users and bundles, the value of user $u$ personalized rating of bundle $b$ can be expressed as follows:
@@
 The sequence of recommendations for bundling recommendations to user $u$ can be obtained by sorting the recommendations according to the personalized score values.
-IV-B4 Model Training
+4.2.4 Model Training
 The model leverages the User Interest Boundary (UIB) loss function, as proposed in the work by Zhuo et al.
@@
 $b_{u}=W^{\text{T}}E_{u}$ (8)
-where $E_{u}$in $R^{d}$ denotes the embedding vector of user $u$ and $W$in $R_{d}$ is the learnable parameter.
+where $E_{u}$in $R^{d}$ denotes the embedding vector of user $u$ and $W$in $R^{d}$ is the learnable parameter.
 The UIB loss is a mixture of point-by-point loss and pairwise loss, which can be formulated as follows:
@@
 Then, we use the AdamW optimizer [26] to minimize the loss function, which can iteratively update the weights of the neural network based on the training data.
-V EXPERIMENTS
-In this section, we delineate the comprehensive evaluation setup for our proposed approach, which we have named ‘HED’.
-The evaluation is meticulously designed to rigorously assess the efficacy, efficiency, and overall performance of ‘HED’ in various scenarios and against multiple benchmarks.
-V-A Experiments Setup
-Datasets.
-We choose the two most common datasets in bundle recommendation from different domains in the real world, Netease [^5] https://music.163.com/ and Youshu [^6] https://www.yousuu.com/, to validate our proposed model.
-Youshu is a book website from China, and the dataset provides interaction data of users, books, and bundles containing a series of books.
-The Netease dataset is constructed by crawling Netease Cloud Music, and it provides interaction data of users, music, and bundles containing a series of music.
-The parameters of the two datasets are shown in Table I.
-TABLE I: Dataset statistics
+Table 1: Dataset statistics
 Dataset NetEase Youshu
@@
 B-I Density 0.06% 0.11%
+5 EXPERIMENTS
+In this section, we delineate the comprehensive evaluation setup for HED.
+The evaluation is meticulously designed to rigorously assess the efficacy, efficiency, and overall performance of HED in various scenarios and against multiple benchmarks.
+5.1 Experimental Setup
+Datasets.
+We choose the two most common datasets in bundle recommendation from different domains in the real world, Netease [^5] https://music.163.com/ and Youshu [^6] https://www.yousuu.com/, to validate our proposed model.
+The Netease dataset is constructed by crawling Netease Cloud Music, and it provides interaction data of users, music, and bundles containing a series of music.
+Youshu is a book website from China, and the dataset provides interaction data of users, books, and bundles containing a series of books.
+The parameters of the two datasets are shown in Table 1.
 Evaluation Metrics.
@@
 Baselines.
-In the subsequent phase of our experiments, we conduct a comprehensive comparison between ‘HED’ and a selection of representative baseline methods.
-This comparative analysis is essential to validate the effectiveness and superior performance of ‘HED’.
+In the subsequent phase of our experiments, we conduct a comprehensive comparison between HED and a selection of representative baseline methods.
+This comparative analysis is essential to validate the effectiveness and superior performance of HED.
 To ensure a fair and thorough comparison, we have chosen baseline methods that are widely recognized and have demonstrated robust performance in similar contexts.
-These methods encompass a range of techniques and algorithms, providing a diverse benchmark for assessing the strengths and potential areas of improvement for ‘HED’.
-Next, we present these baseline methods:
+These methods encompass a range of techniques and algorithms, providing a diverse benchmark for assessing the strengths and potential areas of improvement for HED.
+In specific, we compare HED with the following baseline methods:
   • DAM[9] is a neural network model utilizing attention mechanisms and multi-task learning techniques for bundled recommendations.
@@
   • UHBR[49] is a graph neural network model that unifies multiple associations between users, bundles, and items into a hypergraph and learns using graph convolutional neural networks.
-  • BundleGT[15] is a graph transformer model that contains a token embedding layer, a hierarchical graph transformer layer, and a prediction layer
+  • BundleGT[42] is a graph transformer model that contains a token embedding layer, a hierarchical graph transformer layer, and a prediction layer.
 Training Details.
-In our experimental framework, we allocated 80% of the collected data for training ‘HED’, while the remaining 20% served as the validation set to assess the model’s performance.
+In our experimental framework, we allocated 80% of the collected data for training HED, while the remaining 20% served as the validation set to assess the model’s performance.
 All experiments were carried out on a uniform hardware setup, featuring an Intel(R) Xeon(R) Platinum 8336C CPU and an Nvidia GeForce RTX 4090 GPU, ensuring consistency in computational performance across tests.
 Given the distinct characteristics of our datasets, we tailored the experimental parameters to optimize performance for each.
-Details are listed in Table II.
-TABLE II: The parameter details.
-NetEase Youshu
+A search was conducted to identify an appropriate learning rate within the set range 5e-1, 5e-2, 5e-3, 5e-4, 5e-5, and the rate of 5e-3 was selected for further experimentation.
+Details are listed in Table 2.
+5.2 Comparative Results
+In our comparative analysis, we endeavored to reproduce the baseline methods using publicly available source code or by closely adhering to the methodologies outlined in their respective publications.
+This approach was taken to ensure a fair and accurate comparison with our HED model.
+However, it’s important to note that due to hardware constraints, the method proposed in Wei et al.
+(2023) [42] could not be effectively executed on the Netease dataset.
+Consequently, this limitation has resulted in the absence of data for Recall@80 and NDCG@80 metrics for this particular method.
+Table 2: The parameter details.
+Parameter NetEase Youshu
 Embedding Size 128 64
@@
 Batch Size 2048 1024
+Hypergraph Convolution Dropout 0.1 0.2
 $L$ 2
@@
 Learning Rate 5e-3
-Hypergraph Convolution Dropout 0.2
 U-B graph Convolution Dropout 0.01
-$L_{2}$-norm 0.1
-V-B Comparative Results
-In our comparative analysis, we endeavored to reproduce the baseline methods using publicly available source code or by closely adhering to the methodologies outlined in their respective publications.
-This approach was taken to ensure a fair and accurate comparison with our ‘HED’ model.
-However, it’s important to note that due to hardware constraints, the method proposed in Wei et al.
-(2023) [15] could not be effectively executed on the Netease dataset.
-Consequently, this limitation has resulted in the absence of data for Recall@80 and NDCG@80 metrics for this particular method.
-Table III presents a comprehensive performance comparison between ‘HED’ and the baseline methods.
-TABLE III: Performance comparisons on two real-world datasets with the baselines.
+L2-norm 0.1
+Table 3 presents a comprehensive performance comparison between HED and the baseline methods.
+In the table, the best results for the comparison methods are underlined, while the best progress for HED is shown in bold.
+Additionally, the percentage improvement is calculated.
+The number after HED represents the size of the embeddings.
+The experimental setup is identical to that described in Table 2, with the exception that HED-32 has been trained for 500 epochs on the Youshu dataset.
+Table 3: Performance comparisons on two real-world datasets with the baselines.
 Dataset Modle Recall@20 NDCG@20 Recall@40 NDCG@40 Recall@80 NDCG@80
@@
 UHBR 0.0853 0.0452 0.1335 0.0577 0.1992 0.0732
-BundleGT 0.0903 0.0478 0.1389 0.0606
-HED 0.0963 0.051 0.1448 0.0635 0.212 0.0782
-%Improv. 6.64% 6.69% 4.25% 4.79% 6.04% 6.39%
+BundleGT 0.0903 0.0478 0.1389 0.0606 0.1655 0.0572
+HED-32 0.0895 0.0478 0.1402 0.0608 0.2099 0.0764
+HED-64 0.0934 0.0496 0.1428 0.0624 0.2123 0.0776
+HED-128 0.0963 0.051 0.1459 0.0638 0.213 0.0783
+%Improv. 6.64% 6.69% 5.04% 5.28% 6.93% 6.97%
 Youshu DAM 0.2082 0.1198 0.289 0.1418 0.3915 0.1658
@@
 BundleGT 0.2905 0.1737 0.3909 0.2011 0.5124 0.2242
-HED 0.3153 0.1903 0.423 0.2197 0.5377 0.2452
-%Improv. 3.41% 4.50% 1.54% 3.53% 1.30% 2.90%
+HED-32 0.3157 0.1886 0.4223 0.2175 0.5412 0.2445
+HED-64 0.3165 0.1904 0.4244 0.2197 0.5427 0.2466
+HED-128 0.3122 0.1885 0.421 0.2176 0.5373 0.2438
+%Improv. 3.80% 4.56% 1.87% 3.53% 2.24% 3.48%
 HED has become the new state-of-the-art model for bundled recommendation.
-According to table III, our proposed method is a generalized approach that outperforms the baseline methods of Youshu and NetEase by an average of 2.86% and 5.8%.
+According to Table 3, our proposed method is a generalized approach that outperforms the baseline methods of Youshu and NetEase by an average of 3.25% and 6.26%.
 We attribute this to the complete hypergraph’s superior ability to represent the ternary relationship between users, items, and bundles, and the dual convolution module’s capacity to extract information embedded in this relationship.
@@
 Through the joint analysis of the results of two datasets, we find that the improvements on NetEase are better than those on Youshu.
-The disparity in dataset size and interaction density between the Youshu and NetEase datasets could account for the observed differences.
-Specifically, the Youshu dataset has fewer users, items, and bundles than the NetEase dataset, as well as denser interactions among them, as indicated in Table I.
+The disparity in dataset size and interactions density between the Youshu and NetEase datasets could account for the observed differences.
+Specifically, the Youshu dataset has fewer users, items, and bundles than the NetEase dataset, as well as denser interactions among them, as indicated in Table 1.
 This leads to the model’s $Recall@k$ and $NDCG@k$ metrics showing a significant improvement on the Youshu dataset compared to the NetEase dataset, resulting in a relatively smaller increase in performance on the former.
@@
 The proposed HED is applicable to practical situations.
-The dual convolution module enhances the precision of the package proposition.
-Specifically, although HED performs better than previous models in both $Recall@k$ and $NDCG@k$, the improvement of $NDCG@k$ for the recommendation accuracy measure is greater than that of $Recall@k$, the maximum difference of 1.99%, the minimum difference of 0.05%.
+The dual convolution module enhances the precision of the bundle proposition.
+Specifically, although HED performs better than previous models in both $Recall@k$ and $NDCG@k$, the improvement of $NDCG@k$ for the recommendation accuracy measure is greater than that of $Recall@k$, the maximum difference of 1.66%, the minimum difference of 0.04%.
 This may be due to the fact that previous work does not revert to the recommendation task after extracting information regarding the item.
-In contrast, our dual convolution module, which is designed to mine complete hypergraph information, can enhance the representation of users and bundles based on U-B graph information, thus improving the recommendation accuracy through the final personalized bundle recommendation score Formula 8.
-V-C Ablation Study
+In contrast, our dual convolution module aims to mine the complete hypergraph information, which can enhance the representation of users and bundles based on U-B graph information to improve recommendation accuracy.
+In addition, HED has the highest performance improvement in the Top 20 compared to the Top 40, and the Top 80 also proves the above point.
+5.3 Ablation Study
 In order to better explore the impact of the individual modules of the HED model on the whole, we disassemble it for Ablation Study. We implement four simplified variants of HED:
@@
   • HED--cbu, removing the U-B graph convolution and making both $h_{uu}$ and $h_{bb}$ as zero matrices.
-TABLE IV: Ablated models analysis on Youshu dataset.
-Recall@20 NDCG@20 Recall@40 NDCG@40 Recall@80 NDCG@80
-HED 0.316 (100%) 0.1892 (100%) 0.4209 (100%) 0.2175 (100%) 0.5395 (100%) 0.2445 (100%)
-HED-c 0.3086 (97.66%) 0.1841 (97.3%) 0.4164 (98.93%) 0.2128 (97.84%) 0.5304 (98.31%) 0.2397 (98.04%)
-HED-cu 0.2987 (94.53%) 0.1776 (93.87%) 0.4044 (96.08%) 0.2057 (94.57%) 0.5235 (97.03%) 0.2323 (95.01%)
-HED-cb 0.2965 (93.83%) 0.1779 (94.03%) 0.4025 (95.63%) 0.2064 (94.90%) 0.5267 (97.63%) 0.2337 (95.58%)
-HED-cbu 0.3011 (95.28%) 0.1792 (94.71%) 0.4046 (96.13%) 0.2076 (95.45%) 0.5254 (97.39%) 0.2347 (95.99%)
-We summarize the results of our experiments in Table IV, where the leading number is the performance of the model, and the trailing percentage is the current performance as a percentage of the performance of the full model.
-With the table  IV we can get the following conclusions:
+Table 4: Ablated models analysis on Youshu dataset.
+HED HED-c HED-cu HED-cb HED-cbu
+Recall@20 0.3165 (100%) 0.3086 (96.94%) 0.2987 (94.38%) 0.2965 (93.68%) 0.3011 (95.13%)
+NDCG@20 0.1904 (100%) 0.1841 (96.69%) 0.1776 (93.28%) 0.1779 (93.43%) 0.1792 (94.12%)
+Recall@40 0.4244 (100%) 0.4164 (98.11%) 0.4044 (95.29%) 0.4025 (94.84%) 0.4046 (95.33%)
+NDCG@40 0.2197(100%) 0.2128 (96.86%) 0.2057 (93.63%) 0.2064 (93.95%) 0.2076 (94.49%)
+Recall@80 0.5427 (100%) 0.5304 (97.73%) 0.5235 (96.46%) 0.5267 (97.05%) 0.5254 (96.81%)
+NDCG@80 0.2466 (100%) 0.2397 (97.20%) 0.2323 (94.20%) 0.2337 (94.77%) 0.2347 (95.17%)
+We summarize the results of our experiments in Table 4, where the leading number is the performance of the model, and the trailing percentage is the current performance as a percentage of the performance of the full model.
+With the table  4 we can get the following conclusions:
+Figure 4: Recall and NDCG performance under different $d$
+Figure 5: Recall and NDCG performance under different $n$
+Figure 6: Recall and NDCG performance under different $\alpha$.
   • After excluding the dual convolution module, all six metrics of the model experience a varied decrease, where $NDCG@k$ metrics show a more significant decrease than $Recall@k$ metrics.
@@
 Only one of them leads to a deviation from the complete hypergraph, thus diminishing the model’s performance.
-V-D Sensitivity Analysis
+5.4 Sensitivity Analysis
+Figure 7: Recall and NDCG performance under different $\beta$
 In order to better explore the model’s sensitivity to each hyperparameter, we designed the Sensitivity Analysis experiment.
 Specifically, we experimented with scaling up the hyperparameters in the base experiment by a factor of $2$ or scaling down to $1/2$ of the original and summarized the results in Figures 47.
-Figure 4: Recall and NDCG performance under different $d$
-Figure 5: Recall and NDCG performance under different $n$.
-Figure 6: Recall and NDCG performance under different $\alpha$.
-Figure 7: Recall and NDCG performance under different $\beta$
-  • $d$ represents the dimension, which is the size of the embedding vectors for users, items, and bundles.
-The embedding vectors have a size of 64 for the baseline.
+  • $d$ represents the dimension, which is the size of the embedding for users, items, and bundles.
+The embedding has a size of 64 for the baseline.
 The findings from the sensitivity analysis regarding the parameter $d$ are outlined in Figure 4, the left-hand line graph shows the recall metrics for different $d$ values, and the right-hand line graph shows the NDCG metrics for different $d$ values.
 When $d=16$, there is a more significant decline in model performance than the baseline.
-This is because, when $d$ is too small, the embedding vectors are insufficient in expressing the complete information acquired from the hypergraph and U-B graph.
-On the other hand, when $d>=32$, the embedding vectors exceed the dimensionality limit and can completely express the information derived from the complete hypergraph and U-B graph.
+This is because, when $d$ is too small, the embedding are insufficient in expressing the complete information acquired from the hypergraph and U-B graph.
+On the other hand, when $d>=32$, the embedding exceeds the dimensionality limit and can completely express the information derived from the complete hypergraph and U-B graph.
 The experiments support the claim that for $d>=32$, the model’s performance is similar to that of the baseline.
@@
 […]
-VI Conclusions
+5.5 Discussion on $h_{ii}$ in hypergraphs
+In order to better study our proposed complete hypergraph, we conducted experiments under the Youshu dataset for the case where $h_{ii}$ is the unit matrix.
+The experimental results are shown in Table 5, where although the performance of $Recall@80$ improves, it performs poorly under the more important Top $20$ and Top $40$ items, and there is a significant dip in recommendation accuracy.
+This may be due to the fact that the number of items in the datasets is more than a multiple of the number of users and bundles combined ($2.56$ times in the YouShu dataset, $2.99$ times in the NetEase dataset).
+If $h_{ii}$ is set to a non-zero matrix, it will cause the constructed complete hypergraph to contain too much item information which will affect the quality of the embedding of users and bundles, and affect the final results.
+This is why it performs better than the original model only in the $Recall@80$ metric.
+Table 5: Experiments with $h_{ii}$ on the Youshu dataset.
+$h_{ii}$=E HED
+Recall@20 0.3152 0.3153
+NDCG@20 0.1896 0.1903
+Recall@40 0.4208 0.423
+NDCG@40 0.2184 0.2197
+Recall@80 0.5392 0.5377
+NDCG@80 0.2446 0.2453
+5.6 Implementation efficiency
+The number and efficiency of model parameters are also factors to be considered when evaluating the model.
+For this purpose, we compared the number of parameters of the seven methods on two datasets, and the results are shown in Table 6.
+The majority of the work generated embeddings for users, items, and bundles.
+BundleGT had the highest number of parameters due to the use of multiple graph transformers.
+DAM had the second highest number of parameters due to the incorporation of designing attention for items.
+BGCN, RGNN, and CrossCBR had similar numbers of parameters.
+The MIDGN and UHBR models generate embeddings for users and bundles only.
+MIDGN has a greater number of parameters than UHBR due to the inclusion of the Intent Contrasting module.
+UHBR is the least parameterised model among the existing bundle recommendation algorithms.
+Our approach generates embeddings for users, items, and bundles, so while the number of references is higher than UHBR and MIDGN, it is lower than the other approaches, and our approach can reduce the number of references by adjusting the embedding size with some performance loss.
+Table 6: A comparison of the number of parameters in different models.
+Methods #Params-NetEase #Params-Youshu
+DAM 18490372 5047424
+BGCN 10586112 2941952
+RGNN 10561538 2917378
+MIDGN 4643776 836480
+CrossCBR 10561280 2917120
+UHBR 2649152 819904
+BundleGT 31683840 8751360
+HED-32 5280640 1458560
+HED-64 10561280 2917120
+HED-128 21122560 5834240
+Similarly, we conducted a comparative analysis of the time consumption of the four methods in each period and GPU memory usage on the two datasets.
+The results are presented in Table 7 and Table 8.
+Limited by hardware, the MIGDN method is out of memory on the NetEase dataset.
+The UHBR model employs a smaller hypergraph and graph convolutional neural network, resulting in the fastest computational speed, with a GPU memory usage that is only slightly higher than that of CrossCBR.
+The CrossCBR method employs the lightGCN algorithm and a comparative learning framework with the lowest GPU memory usage, and is second only to UHBR in computational speed among the four methods.
+The MIDGN approach necessitates the generation of additional adjacency matrices during the graph decomposition process, which requires a significant amount of GPU memory and also results in a slower computational speed.
+BundleGT, on the other hand, runs slower and has a high GPU memory usage because of the graph transformer.
+Our method, on the other hand, exhibits different results depending on the dataset.
+On the Youshu dataset, our method uses a moderate amount of GPU memory and runs fast with the highest performance.
+With regard to the more complex NetEase dataset, it should be noted that the optimal results achieved were at the cost of high GPU memory usage and long training times.
+Furthermore, it is important to highlight that this level of enhancement was higher than that observed in the Youshu dataset.
+This is a result of the deeper mining of data, which included the construction of larger complete hypergraphs and the utilisation of dual convolutional modules in order to improve the accuracy of the recommendations.
+In addition, adjusting the embedding size only affects the model training time and does not reduce the GPU memory usage.
+Due to the limitations of our existing hardware, HED is out of memory on the iFashion dataset, and the application of our method to the industrial sector will also suffer from such problems.
+The recommender system utilised in industry comprises four distinct phases: recall [3], ranking [34] and re-ranking [29].
+The deployment of our method in the recall phase necessitates a significant amount of memory space, whereas the deployment in the other phases circumvents the issue of hardware requirements.
+Table 7: Implementation costs of different methods on NetEase datase.
+Methods Training Inference GPU memory
+MIDGN OOM OOM OOM
+CrossCBR 5 2 2728MiB
+UHBR 9 3 8412MiB
+BundleGT 180.07586 177.97326 10432MiB
+HED-32 27 2 23148MiB
+HED-64 36 2 23148MiB
+HED-128 56 4 23148MiB
+Table 8: Implementation costs of different methods on Youshu datase.
+Methods Training Inference GPU memory
+MIDGN 7 8 5028MiB
+CrossCBR 2.4 2 1642MiB
+UHBR 0.71 1 2388MiB
+BundleGT 29.1437 15 5392MiB
+HED-32 1.69 1 3232MiB
+HED-64 1.86 1 3232MiB
+HED-128 2.32 1 3232MiB
+6 Conclusions
 In this study, the significance of creating user-item-bundle triples in bundle recommendation is examined.
-A comprehensive hypergraph construction approach and HED model are introduced.
-A comprehensive hypergraph is a technique for conveying the ternary relationship formed by user, item, and bundle interactions.
+This paper introduces a complete hypergraph construction method and hypergraph-enhanced dual convolutional neural network (HED) models that are capable of utilising the constructed complete hypergraph.
+Comprehensive hypergraphs represent a methodology for elucidating the ternary relationships that emerge from user, item, and bundle interactions.
+This approach considers the interactions between the three entities, as well as the relationships between users and between bundles, which include interactions between each entity and itself.
 The HED model includes an embedding initialization layer, a dual convolutional layer, and a prediction layer.
-The embedding initialization layer produces embedding vectors for every user, item, and bundle.
-The dual convolutional layer extracts complete hypergraph information and enhances the embedding vectors by merging the U-B graph information.
-The prediction layer provides recommendations based on the final embedding vectors.
-A substantial volume of experiments conducted on Youshu and NetEase empirically prove the superiority of the HED model we have proposed.
+The embedding initialization layer produces embedding for every user, item, and bundle.
+The dual convolutional layers extract the complete hypergraph information and enhance the embedding by fusing the most important U-B graph information.
+The prediction layer provides recommendations based on the final embedding.
+Comparative experiments on Youshu and NetEase datasets demonstrate that our proposed HED model represents a new state-of-the-art approach to the field of bundle recommendation.
+It is notable that the HED outperforms the Youshu dataset on the more complex NetEase dataset, and that it significantly improves recommendation accuracy on both datasets.
+HED effectiveness was demonstrated by performing ablation studies and sensitivity analyses revealing the working mechanism.
+In addition, we discuss the impact of hypergraph construction on the results, analysing the efficiency of the model implementation in terms of parametric runtime and GPU memory usage.
+Although HED is the recommended SOTA model for the current bundle, it is important to note that HED still has certain shortcomings.
+The construction of complete hypergraphs necessitates a significant allocation of memory, which consequently increases the hardware requirements for deployment of the model.
+Further research is required to identify methods of reducing the memory footprint and improving the learning efficiency.
+Furthermore, this representation of ternary relations and the mechanism of learning and enhancing the representation of ternary relations through the most important binary relations can be generalised to other problems, such as group recommendation in recommender systems.
 Acknowledgment
-This research is supported by the Fundamental Research Funds for the Central Universities (3102021ZDHQD09).
+This research is supported by the National Natural Science Foundation of China (No.62103330) and Guangdong Basic and Applied Basic Research Foundation (2024A1515012388).
+[…]